# Generic motives and motivic cohomology of fields

### v1

@@ 5.1. Borel’s computations and $\lambda$-structures
As it is very involved, we indicate how to deduce the proof of the above theorem from [Bor74], Th. 7.5.
One considers the maximal compact open subgroup $K_{r}=\operatorname{SO}_{r}^{r_{1}}\times\operatorname{SU}_{r}^{r_{2}}$ of $G_{r}$ and the associated symmetric space $X_{r}=K_{r}\backslash G_{r}$. The group $\Gamma_{r}$ acts properly discontinuously and almost freely on $X_{r}$ so that the topological quotient $O_{r}=X_{r}/\Gamma_{r}$ is an orbifold (see e.g. [Thu22]). The canonical map $\operatorname{B}\Gamma_{r}=X_{r}\mathbin{/\mkern-6.0mu/}G_{r}\rightarrow O_{r}$ is a $\mathbb{Q}$-equivalence (7.0.1). [^34] According to Selberg’s lemma, there existe a finite index normal torsion-free subgroup $\Gamma^{\prime}_{r}\subset\Gamma_{r}$. Then $X_{r}/\Gamma^{\prime}_{r}$ is a manifold and a model for $\operatorname{B}\Gamma_{r}$. But now the canonical map $X_{r}/\Gamma^{\prime}_{r}\rightarrow O_{r}$ is a finite branched cover, and therefore a $\mathbb{Q}$-equivalence. The de Rham isomorphism for orbifolds (see [Sat56]) implies that one can compute the singular cohomology of $O_{r}$ with real coefficients, and therefore that of $\operatorname{B}\Gamma_{r}$, as the cohomology of the differential graded $\mathbb{R}$-algebra $\Omega^{*}(O_{r})\simeq\Omega^{*}(X_{r})^{\Gamma_{r}}$. On the other hand, according to the (topological version) of the Poincaré lemma, the continuous cohomology of $G_{r}$ can be identified with the cohomology of the $\mathbb{R}$-dga $\Omega^{*}(X_{r})^{G_{r}}$. The Mastushima map considered in [Bor74, §3.1] is nothing else than the inclusion map of real differential graded algebras
@@
The latter is therefore a quasi-isomorphism in degree less than $r/4$ according to the conjunction of Th. 7.5 and §9.5, formula (3) of [Bor74]. Finally, it is straightforward to check that $j_{\Gamma_{r}}$ models the restriction map $\nu_{r}^{*}$ on continuous cohomology. ∎
@@
Indeed, for any $n\leq r/4$, $H^{n}_{ct}(G_{r},\mathbb{R})$ has finite dimension over $\mathbb{R}$, and does not depend on $r$ [^35] This follows from the van Est isomorphism, which allows to identify the latter cohomology groups with the singular cohomology group $H^{n}_{sing}(G_{r}^{u},\mathbb{R})$ of the compact real form $G_{r}^{u}$ associated with the semi-simple real Lie group $G_{r}$. See below. It follows that both the source and the target of $\nu_{r}^{*}$ satisfies the Mittag-Leffler condition, and this implies that $\nu^{*}$ can be identified with the projective limit of the $\nu_{r}^{*}$. Thus the preceding theorem allows us to conclude.
@@
Recall that according to the universal coefficient theorem, one always gets a canonical isomorphism:
@@
where $\vee$ means the dual of (discrete) $\mathbb{R}$-vector spaces. Therefore, as this $\mathbb{R}$-vector space is finite dimensional according to the preceding corollary, one deduces a canonical isomorphism
@@
We will define the following $\mathbb{R}$-valued continuous homology [^36] Because of finite dimensionality, it coincides with other possible definition of continuous homology via derived functors.:
@@
Consider the above notation. Then for any integer $n\geq 2$, the Hurewicz map composed with the above isomorphism $\nu_{*}$ induces an isomorphism of $\mathbb{R}$-vector spaces:
@@
This is a nice formula, which already allows to recover the rank of K-theory. However, we do not know how to determine the gamma filtration on the right hand-side directly, though it exists naturally (see below). Therefore we have to come back to the known strategy and use the compact form of the ind-Lie group $\operatorname{SL}(K\otimes_{\mathbb{Q}}\mathbb{R})$.
@@
We now recall the necessary computation from Lie groups and Lie algebras. To the ind-Lie group $G=\operatorname{SL}(K\otimes_{\mathbb{Q}}\mathbb{R})=\operatorname{SL}(\mathbb{R})^{r_{1}}\times\operatorname{SL}(\mathbb{C})^{r_{2}}$, one associates using Cartan duality the following objects:
@@
$\varphi:H^{*}_{ct}(G,\mathbb{R})\underset{(1234)}{\xrightarrow{\ \sim\ }}H^{*}_{dR}(X^{u})\xrightarrow{\sim}H^{*}_{sing}(X^{u},\mathbb{R}).$
@@
$\rho_{B}:K_{n}(\mathcal{O}_{K})\otimes_{\mathbb{Q}}\mathbb{R}\xrightarrow{\nu_{*}}\operatorname{P}_{n}H_{*}^{ct}\big{(}\operatorname{SL}(K\otimes_{\mathbb{Q}}\mathbb{R}),\mathbb{R}\big{)}\xrightarrow{\varphi^{\vee}}\operatorname{P}_{n}H_{*}^{sing}\big{(}(\operatorname{SU}/\operatorname{SO})^{r_{1}}\times\operatorname{SU}^{r_{2}},\mathbb{R}\big{)}$
called the Borel regulator associated with the number field $K$ (in degree $n$).
@@
    Beware that Burgos has explained with a remarkable precision in [BG02, §9.5] that one needs to renormalize the above regulator in order to compare it with the Beilinson regulator (up to multiplying by $2$ on the factors corresponding to the $r_{1}$ real places). In our approach, this problem is irrelevant.
@@
    The only non-canonical isomorphism that composes the Borel regulator is the isomorphism (2) mentioned above: it is noted, however, that it is degree-wise a multiplication by a scalar within the ambient complex ind-Lie algebra $\mathfrak{g}_{\mathbb{C}}$.
@@
We are now in position to elucidate the behavior of the Borel regulator with respect to the canonical $\lambda$-structure on the left hand-side, and an appropriate computable $\lambda$-structure on the right hand-side.
To define the $\lambda$-operations, we use the method of Soulé in [Sou85, §1]. In our case, this leads to consider the exterior power of matrices. After taking the obvious base on $\Lambda^{r}(\mathbb{Z}^{n})$, this gives a morphism of schemes
$\Lambda^{r}_{n}:\operatorname{SL}_{n}\rightarrow\operatorname{SL}_{\binom{n}{k}}$
and as $n$ goes to $\infty$, a morphism of ind-group schemes $\Lambda^{r}:\operatorname{SL}\rightarrow\operatorname{SL}$. According to the construction of loc. cit., this morphism evaluated at $\mathcal{O}_{K}$ induces by functoriality a morphism on $K_{n}(\mathcal{O}_{K})=\pi_{n}(\operatorname{BSL}(\mathcal{O}_{K})^{+})$ which is nothing else than the operation $\lambda^{k}$ defined in loc. cit.
By the same formula, we obtain operations $\Lambda^{k}$ on $\operatorname{SU}$ and $\operatorname{SO}$ respectively, and therefore by functoriality an operation on the target of the Borel regulator $\rho_{B}$. It follows from the construction, as we have detailed, that $\rho_{B}$ commutes with the action just constructed on both its source and target.
It remains to determine the induced structure on $H_{*}^{sing}\big{(}(\operatorname{SU}/\operatorname{SO})^{r_{1}}\times\operatorname{SU}^{r_{2}},\mathbb{R}\big{)}$ and on its primitive part.
For this, we recall the theory of characteristic classes for notation. Given a topological group $\mathfrak{G}$, we know that there exists universal principal topological bundle $E\mathfrak{G}$ classifying respectively the special (i.e. with trivial determinant) complex and real vector bundles (see e.g. [Mil56]). One deduces a fibration sequence:
@@
$H^{n+1}_{sing}(\operatorname{B}\mathfrak{G},\mathbb{R})\xrightarrow{s^{*}}H^{n+1}_{sing}(\Sigma\mathfrak{G},\mathbb{R})\simeq H^{n}_{sing}(\mathfrak{G},\mathbb{R}).$
Note that this map is natural with respect to the topological group $\mathfrak{G}$. In our case, where $\mathfrak{G}=\operatorname{SU},\operatorname{SO}$ is an ind-Lie group, we deduce from the Serre spectral sequence associated with the fibration $p$ that the suspension map is an isomorphism. By taking real duals, one deduces the following homological suspension isomorphisms which fit into the commutative diagram:
@@
where $\iota:\operatorname{SO}\rightarrow\operatorname{SU}$ is the natural inclusion. Now, the theory of characteristic classes (see e.g. [Swi02, Chap. 16]) applies and gives:
$H^{*}_{sing}(\operatorname{B}\operatorname{SU},\mathbb{R})$ $=\mathbb{R}[c_{2},\ldots,c_{i},\ldots]$
$H^{*}_{sing}(\operatorname{B}\operatorname{SO},\mathbb{R})$ $=\mathbb{R}[p_{2},\ldots,p_{j},\ldots]$
@@
$\iota^{*}(c_{i})=\begin{cases}0&i\text{ odd,}\\ (-1)^{j}(2j-1)!.p_{j}+\text{lower degrees}&i=2j.\end{cases}$
@@
The real singular homology with its weight graduation coming from the preceding $\lambda$-structure
$H_{*}^{sing}\big{(}(\operatorname{SU}/\operatorname{SO})^{r_{1}}\times\operatorname{SU}^{r_{2}},\mathbb{R}\big{)}$
is the free exterior real algebra generated by $r_{2}$ classes of degree $12i-$ 1 and weight $i$ for any $i\geq 2$ and by $r_{1}$ classes of degree $12j-$ 1 and weight $j$ for odd $j\geq 2$.
This follows from the preceding discussion and the (dual) Künneth formula. Let us draw the explicit corollary, which achieves the proof of 5.1.4.
@@
Given a number field $K$ with $r_{1}$, $r_{2}$ real and complex places respectively, a basis for the K-theory graded $\mathbb{Q}$-vector space $K_{*}(\mathcal{O}_{K})_{\mathbb{Q}}$ restricted to degrees 2$\geq$ 2∗ ≥ 2 is given by:
  • $r_{2}$ elements in degree $12i-$ 1 and weight $i$ for $i\geq 2$,
  • $r_{1}$ elements in degree $12j-$ 1 and weight $j$ for odd $j\geq 2$.
@@ 5.2. The effect of transcendental elements
We will now deduce from the computations of generic motives obtained in the preceding section some information about motivic cohomology of fields. All the next results are based on formula (5.0.1.a). First, as a corollary of 4.1.2 and 5.1.4, we get the following computation:
@@
As there are infinitely many points $x\in\mathbb{A}^{1}_{K,(0)}$, whose residue field $\kappa_{x}$ has at least one complex place, one deduces from 5.1.4 that for all $i\geq 2$, the abelian group $\operatorname{H}_{\mathcyr{B}}^{n,i}(K(t))$ has infinite rank.
This picture can be confirmed in higher transcendence degree. Indeed, applying now 4.3.1, one obtains the following proposition.
@@
Let $K$ be a field of characteristic 00 such that $\operatorname{degtr}K\geq d>0$.
@@
Indeed, by assumptions, $K$ contains the field of rational functions $\mathbb{Q}(t_{1},...,t_{d})$. According to 4.1.4, one can assume that $K=\mathbb{Q}(t_{1},,...,t_{d})$. According to 4.3.1, one deduces that $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$ contains $\bigoplus_{x\in\mathbb{A}^{1}_{\mathbb{Q},(0)}}\operatorname{H}_{\mathcyr{B}}^{1,i-n+1}(\kappa_{x})$ which is infinite according to 5.1.4 and the fact there exists an infinity of finite monogeneous extensions of $\mathbb{Q}$ with at least on complex place. ∎
@@
According to the splitting $\varphi$ given in 4.1.2, it is easy to construct infinite free families of elements of $\operatorname{H}_{\mathcyr{B}}^{n,i}(\mathbb{Q}(t_{1},...,t_{n}))$ under the assumptions of the preceding proposition. As a function field $K/k$ of transcendence degree $d=n$ is a finite extension of $\mathbb{Q}(t_{1},...,t_{n})$ one deduces infinite free families of $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$. We let these constructions as an exercise to the reader. [^39] As a hint, consider product of symbols extracted from some transcendental basis of $K/\mathbb{Q}$, elements coming from the Borel regulator on an appropriate number field $E$ and eventually restriction or corestriction along finite field extensions on motivic cohomology. All these operations can be realized geometrically as well by taking appropriate smooth models.
One can even push this argument further to get non-finiteness results of motivic cohomology.
@@
Let $K$ be an extension field of $\mathbb{Q}$ with infinite transcendence degree. Then for any integers $i\geq n\geq 2$, the abelian group $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$ has rank equal to $\operatorname{card}(K)$.
An important remark is that this result extends an earlier theorem due to Springer, which concerns the Milnor K-theory part $n=i$: see [BT73, 5.10] (and below for more discussion).
@@
We next consider the other inequality. By assumption, there exists a subfield $F\subset K$ and a finite morphism $F(t_{1},...,t_{n})\rightarrow K$. Using 4.1.4, it suffices to treat the case of $K=F(t_{1},...,t_{n})$. Then according to 4.3.1, one deduces that $\operatorname{H}_{\mathcyr{B}}{n,i}(K)$ contains as a direct factor the abelian group $\oplus_{x\in\mathbb{A}^{1}_{F,{(0)}}}\operatorname{H}_{\mathcyr{B}}^{1,i-n+1}(\kappa_{x})$. Applying again 4.1.4, the latter groups contains for example $\oplus_{x\in F}\operatorname{H}_{\mathcyr{B}}^{1,i-n+1}(\mathbb{Q}[i])$ whose rank is bigger or equal to $\operatorname{card}(F)=\operatorname{card}(K)$. ∎
@@
As a conclusion, there is a range, controlled by the transcendence degree over $\mathbb{Q}$, in which all motivic cohomology groups of fields are infinite. However, a tempting conjecture emerges that we will now formulate. We will directly state it in arbitrary characteristics though we have focused on the characteristic 00 case. Recall one defines the Kronecker dimension of a field $K$ as the integer
$\delta(K)=\begin{cases}\operatorname{degtr}(K/\mathbb{Q})+1&\operatorname{char}(K)=0\\ \operatorname{degtr}(K/\mathbb{F}_{p})&\operatorname{char}(K)=p>0.\end{cases}$
Conjecture.
@@
unless $(n,i)=(0,0)$ or $n\in[1,\delta(K)]$.
There is few evidences for this conjecture. Nevertheless, it holds for the fields $K$ and $K(t)$ such that $\delta(K)\leq 1$: as we have seen, the case of number fields follows from 5.1.4 and 5.2.1. The case of characteristic $p>0$ is a consequence of [Qui72] and [Har77] (see below). We will end-up this section with some further evidences.
5.2.6.
The Beilinson-Soulé conjecture. Actually the above conjecture in the range $n<1$ is already well-known as the Beilinson-Soulé conjecture, in its “strong” form. This terminology appears in the introduction of [Lev93]. By opposition, the “weak” form of the Beilinson-Soulé conjecture asks for the vanishing of motivic cohomology in degree $<0$ (see e.g. [SV00, §3, Be2]). [^40] The integral version is equivalent to the rational one: see [Kah05, Lem. 24]. We will simply say Beilinson-Soulé conjecture in the sequel for the strong form.

### v3

@@ 5.1. Borel’s computations and $\lambda$-structures
As the proof is very involved, we indicate how to deduce the proof of the above theorem from [Bor74], Theorem 7.5.
One considers the maximal compact open subgroup $K_{r}=\operatorname{SO}_{r}^{r_{1}}\times\operatorname{SU}_{r}^{r_{2}}$ of $G_{r}$ and the associated symmetric space $X_{r}=K_{r}\backslash G_{r}$. The group $\Gamma_{r}$ acts properly discontinuously and almost freely on $X_{r}$ so that the topological quotient space $O_{r}=X_{r}/\Gamma_{r}$ is an orbifold (see e.g. [Thu22, Chap. 13]). Moreover, almost by construction of an orbifold, the natural map $\operatorname{B}\Gamma_{r}:=X_{r}\mathbin{/\mkern-6.0mu/}\Gamma_{r}\rightarrow O_{r}$ between the homotopy quotient to the orbifold quotient is a $\mathbb{Q}$-equivalence, in the sense of 6.3.6.424242In the original argument given by Borel, one applies Selberg’s lemma (see [Alp87]) to get a finite index normal torsion-free subgroup $\Gamma^{\prime}_{r}\subset\Gamma_{r}$. Then $\Gamma^{\prime}_{r}$ acts freely on, $X_{r}$ so that $X_{r}/\Gamma^{\prime}_{r}$ is a manifold and a model for $\operatorname{B}\Gamma^{\prime}_{r}$. But now the canonical map $X_{r}/\Gamma^{\prime}_{r}\rightarrow X_{r}/\Gamma_{r}=O_{r}$ is a finite branched cover, and therefore a $\mathbb{Q}$-equivalence. Moreover, the canonical map $B\Gamma^{\prime}_{r}\rightarrow B\Gamma_{r}$ is also a $\mathbb{Q}$-equivalence (see e.g. [Bro94, Chap. III.7]). One deduces by composition a natural $\mathbb{Q}$-equivalence $B\Gamma_{r}\simeq O_{r}$.
The de Rham isomorphism theorem for orbifolds (see [Sat56]) implies that one can compute the singular cohomology of $O_{r}$ with real coefficients, and therefore that of $\operatorname{B}\Gamma_{r}$, as the cohomology of the differential graded $\mathbb{R}$-algebra $\Omega^{*}(O_{r})\simeq\Omega^{*}(X_{r})^{\Gamma_{r}}$. On the other hand, according to the topological version of the Poincaré lemma, the continuous cohomology of $G_{r}$ can be identified with the cohomology of the $\mathbb{R}$-dga $\Omega^{*}(X_{r})^{G_{r}}$. The Matsushima map considered in [Bor74, §3.1] is nothing else than the inclusion map of real differential graded algebras
@@
The latter is therefore a quasi-isomorphism in degree less than $r/4$ according to the conjunction of Th. 7.5 and §9.5, formula (3) of [Bor74]. Finally, it is straightforward to verify that $j_{\Gamma_{r}}$ models the restriction map $\nu_{r}^{*}$ on continuous cohomology. ∎
@@
Indeed, for any $n\leq r/4$, $H^{n}_{ct}(G_{r},\mathbb{R})$ has finite dimension over $\mathbb{R}$, and does not depend on $r$ [^43] This follows from the van Est isomorphism, which allows to identify the latter cohomology groups with the singular cohomology group $H^{n}_{\text{sing}}(G_{r}^{u},\mathbb{R})$ of the compact real form $G_{r}^{u}$ associated with the semi-simple real Lie group $G_{r}$. See below. It follows that both the source and the target of $\nu_{r}^{*}$ satisfy the Mittag-Leffler condition, implying that $\nu^{*}$ can be identified with the projective limit of the $\nu_{r}^{*}$. Thus the preceding theorem allows us to conclude.
@@
Recall that according to the universal coefficient theorem, one gets a canonical isomorphism:
@@
where $\vee$ denotes the dual of $\mathbb{R}$-vector spaces. Since this $\mathbb{R}$-vector space is finite dimensional according to the preceding corollary, one deduces a canonical isomorphism
@@
We will define the following $\mathbb{R}$-valued continuous homology [^44] Due to finite dimensionality, it coincides with the other possible definition of continuous homology via derived functors.:
@@
Consider the above notation. Then for any integer $n\geq 2$, the Hurewicz map, composed with the above isomorphism $\nu_{*}$, induces an isomorphism of $\mathbb{R}$-vector spaces:
@@
This is a nice formula, which already allows us to recover the rank of K-theory. However, we do not know how to determine the gamma filtration on the right-hand side directly, though it exists naturally (see below). Therefore, we must return to the known strategy and use the compact form of the ind-Lie group $\operatorname{SL}(K\otimes_{\mathbb{Q}}\mathbb{R})$.
@@
We now recall the necessary computations from Lie groups and Lie algebras. With the ind-Lie group $G=\operatorname{SL}(K\otimes_{\mathbb{Q}}\mathbb{R})=\operatorname{SL}(\mathbb{R})^{r_{1}}\times\operatorname{SL}(\mathbb{C})^{r_{2}}$, one associates using Cartan duality the following objects:
@@
$\varphi:H^{*}_{ct}(G,\mathbb{R})\underset{(1234)}{\xrightarrow{\ \sim\ }}H^{*}_{dR}(X^{u})\xrightarrow{\sim}H^{*}_{\text{sing}}(X^{u},\mathbb{R}).$
@@
$\nu_{\text{Bo}}:K_{n}(\mathcal{O}_{K})\otimes_{\mathbb{Z}}\mathbb{R}\xrightarrow{\nu_{*}}\operatorname{P}_{n}H_{*}^{ct}\big{(}\operatorname{SL}(K\otimes_{\mathbb{Q}}\mathbb{R}),\mathbb{R}\big{)}\xrightarrow{\varphi^{\vee}}\operatorname{P}_{n}H_{*}^{\text{sing}}\big{(}(\operatorname{SU}/\operatorname{SO})^{r_{1}}\times\operatorname{SU}^{r_{2}},\mathbb{R}\big{)}$
called the Borel isomorphism associated with the number field $K$ (in degree $n$).
Note that the Borel regulator is then obtained by the following composite map:
(5.1.9.a) $\rho_{\text{Bo}}:K_{n}(\mathcal{O}_{K})_{\mathbb{Q}}\rightarrow K_{n}(\mathcal{O}_{K})\otimes_{\mathbb{Z}}\mathbb{R}\xrightarrow{\nu_{\text{Bo}}}\operatorname{P}_{n}H_{*}^{\text{sing}}\big{(}(\operatorname{SU}/\operatorname{SO})^{r_{1}}\times\operatorname{SU}^{r_{2}},\mathbb{R}\big{)}.$
@@
    Beware that Burgos has explained with remarkable precision in [BG02, §9.5] that one needs to renormalize the above regulator to compare it with the Beilinson regulator (up to multiplying by $2$ on the factors corresponding to the $r_{1}$ real places). In our approach, this problem is irrelevant.
@@
    The only non-canonical isomorphism involved in the Borel regulator is the isomorphism (2) mentioned above. However, it is noted that it is degree-wise a multiplication by a scalar within the complexified ind-Lie algebra $\mathfrak{g}_{\mathbb{C}}$.
@@
We are now in a position to elucidate the behavior of the Borel regulator with respect to the canonical $\lambda$-structure on the left-hand side, and an appropriate computable $\lambda$-structure on the right-hand side.
To define the $\lambda$-operations, we use the method of Soulé in [Sou85, §1]. In our case, this leads us to consider the exterior power of matrices. After taking the obvious base on $\Lambda^{r}(\mathbb{Z}^{n})$, we deduce a morphism of schemes
$\Lambda^{r}_{n}:\operatorname{SL}_{n}\rightarrow\operatorname{SL}_{\binom{n}{r}}$
and as $n$ goes to $\infty$, a morphism of ind-group schemes $\Lambda^{r}:\operatorname{SL}\rightarrow\operatorname{SL}$. According to the construction of loc. cit., this morphism evaluated at $\mathcal{O}_{K}$ induces by functoriality a morphism on $K_{n}(\mathcal{O}_{K})=\pi_{n}(\operatorname{BSL}(\mathcal{O}_{K})^{+})$ which is nothing else than the operation $\lambda^{r}$ defined in loc. cit.
By the same formula, we obtain operations $\Lambda^{r}$ on $\operatorname{SU}$ and $\operatorname{SO}$ respectively, and therefore by functoriality an operation on the target of the Borel regulator $\rho_{B}$. It follows from the construction, as we have detailed, that $\rho_{B}$ commutes with the action just constructed on both its source and target.
It remains to determine the induced structure on $H_{*}^{\text{sing}}\big{(}(\operatorname{SU}/\operatorname{SO})^{r_{1}}\times\operatorname{SU}^{r_{2}},\mathbb{R}\big{)}$ and on its primitive part.
For this, we recall the theory of characteristic classes for notation. Given a topological group $\mathfrak{G}$, there exists a universal principal topological bundle $E\mathfrak{G}$ classifying respectively the special464646i.e. with trivial determinant complex and real vector bundles (see e.g. [Mil56]). One deduces the classical fibration sequence:
@@
$H^{n+1}_{\text{sing}}(\operatorname{B}\mathfrak{G},\mathbb{R})\xrightarrow{s^{*}}H^{n+1}_{\text{sing}}(\Sigma\mathfrak{G},\mathbb{R})\simeq H^{n}_{\text{sing}}(\mathfrak{G},\mathbb{R}).$
Note that this map is natural with respect to the topological group $\mathfrak{G}$. In our case, where $\mathfrak{G}=\operatorname{SU},\operatorname{SO}$ is an ind-Lie group, we deduce from the Serre spectral sequence associated with the fibration $p$, and the known cohomologies of $\mathfrak{G}$ and $B\mathfrak{G}$ (see e.g. [Swi02, Chap. 16]), that the suspension map is an isomorphism. By taking real duals, one deduces the following homological suspension isomorphisms which fit into the commutative diagram:
@@
where $\iota:\operatorname{SO}\rightarrow\operatorname{SU}$ is the natural inclusion. Now, the theory of characteristic classes (see again [Swi02, Chap. 16]) applies and gives:
$H^{*}_{\text{sing}}(\operatorname{B}\operatorname{SU},\mathbb{R})$ $=\mathbb{R}[c_{2},\ldots,c_{i},\ldots]$
$H^{*}_{\text{sing}}(\operatorname{B}\operatorname{SO},\mathbb{R})$ $=\mathbb{R}[p_{2},\ldots,p_{j},\ldots]$
@@
$\iota^{*}(c_{i})=\begin{cases}0&i\text{ odd,}\\ (-1)^{j}(2j-1)!p_{j}+\text{lower degrees}&i=2j.\end{cases}$
@@
The real singular homology with its weight graduation coming from the previously defined $\lambda$-structure
$H_{*}^{\text{sing}}\big{(}(\operatorname{SU}/\operatorname{SO})^{r_{1}}\times\operatorname{SU}^{r_{2}},\mathbb{R}\big{)}$
is the free exterior real algebra generated by $r_{2}$ classes of degree $12i-$ 1 and weight $i$ for each $i\geq 2$ and by $r_{1}$ classes of degree $12j-$ 1 and weight $j$ for each odd $j\geq 2$.
This follows from the preceding discussion and the (dual) Künneth formula. Let us draw the explicit corollary, which completes the proof of 5.1.4.
@@
Given a number field $K$ with $r_{1}$ real and $r_{2}$ complex places, a basis for the K-theory graded $\mathbb{Q}$-vector space $K_{*}(\mathcal{O}_{K})_{\mathbb{Q}}$ restricted to degrees 2$\geq$ 2∗ ≥ 2 is given by:
  • $r_{2}$ elements in degree $12i-$ 1 and weight $i$ for each $i\geq 2$,
  • $r_{1}$ elements in degree $12j-$ 1 and weight $j$ for each odd $j\geq 2$.
@@ 5.2. The effect of transcendental elements
We will now deduce from the computations of generic motives obtained in the preceding section some information about motivic cohomology of fields. All subsequent results are based on formula (5.0.1.a). First, as a corollary of 4.2.2 and 5.1.4, we get the following computation:
@@
As there are infinitely many points $x\in\mathbb{A}^{1}_{K,(0)}$ whose residue field $\kappa_{x}$ has at least one complex place, one deduces from 5.1.4 that for all $i\geq 2$, the abelian group $\operatorname{H}_{\mathcyr{B}}^{n,i}(K(t))$ has infinite rank.
This picture can be confirmed in higher transcendence degree. Indeed, applying now 4.4.1, one obtains the following proposition.
@@
Let $K$ be a field of characteristic 00 such that $\operatorname{trdeg}K\geq d>0$.
@@
Indeed, by assumption, $K$ contains the field of rational functions $\mathbb{Q}(t_{1},...,t_{d})$. According to 4.2.4, one can assume that $K=\mathbb{Q}(t_{1},,...,t_{d})$. According to 4.4.1, it follows that $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$ contains $\bigoplus_{x\in\mathbb{A}^{1}_{\mathbb{Q},(0)}}\operatorname{H}_{\mathcyr{B}}^{1,i-n+1}(\kappa_{x})$ which is infinite according to 5.1.4 and the fact there exists infinity many finite monogeneous extensions of $\mathbb{Q}$ with at least one complex place. ∎
@@
According to the splitting $\varphi$ given in 4.2.2, it is straightforward to construct infinitely many free families of elements of $\operatorname{H}_{\mathcyr{B}}^{n,i}(\mathbb{Q}(t_{1},...,t_{n}))$ under the assumptions of the previous proposition. As a function field $K/k$ of transcendence degree $d=n$ is a finite extension of $\mathbb{Q}(t_{1},...,t_{n})$ one deduces infinitely many free families of $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$. We leave these constructions as an exercise to the reader. [^48] As a hint, consider products of symbols extracted from some transcendental basis of $K/\mathbb{Q}$, elements coming from the Borel regulator on an appropriate number field $E$ and eventually applying restriction or corestriction along finite field extensions on motivic cohomology. All these operations can be realized geometrically as well by taking appropriate smooth models.
One can further extend this argument to obtain non-finiteness results of motivic cohomology.
@@
Let $K$ be an extension field of $\mathbb{Q}$ of infinite transcendence degree. Then for any pair of integers $(n,i)$ with $i\geq n\geq 2$, the abelian group $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$ has rank equal to $\operatorname{card}(K)$.
It is noteworthy that this result extends an earlier theorem by Springer, concerning the Milnor K-theory case (i.e. $n=i$): see [BT73, 5.10] (and below for more discussion).
@@
Next, we consider the opposite inequality. By assumption, there exists a subfield $F\subset K$ and a finite morphism $F(t_{1},\ldots,t_{n})\rightarrow K$. Applying 4.2.4, it suffices to treat the case of $K=F(t_{1},\ldots,t_{n})$. According to 4.4.1, it follows that $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)$ contains, as a direct factor, the abelian group
$\oplus_{x\in\mathbb{A}^{1}_{F,{(0)}}}\operatorname{H}_{\mathcyr{B}}^{1,i-n+1}(\kappa_{x}).$
Applying again 4.2.4, the latter groups contains for instance the abelian group
$\oplus_{x\in F}\operatorname{H}_{\mathcyr{B}}^{1,i-n+1}(\mathbb{Q}[i])$
whose rank is bigger or equal to $\operatorname{card}(F)=\operatorname{card}(K)$. ∎
@@
As a conclusion, there is a range, controlled by the transcendence degree over $\mathbb{Q}$, in which all motivic cohomology groups of fields have infinite rank. Nevertheless, a compelling conjecture arises, which we now state. We will state it directly in arbitrary characteristic, even though our focus has been on the characteristic 00 case. Recall that the Kronecker dimension of a field $K$ is given by
$\delta(K)=\begin{cases}\operatorname{trdeg}(K/\mathbb{Q})+1&\operatorname{char}(K)=0\\ \operatorname{trdeg}(K/\mathbb{F}_{p})&\operatorname{char}(K)=p>0.\end{cases}$
Conjecture 5.2.6.
@@
unless $n=i=0$ or $n\in[1,\delta(K)]$.
There is little evidence for this conjecture. Nevertheless, it holds for fields $K$ and $K(t)$ such that $\delta(K)\leq 1$: as we have seen, the case of number fields follows from 5.1.4 and 5.2.1. The case of characteristic $p>0$ is a consequence of [Qui72] and [Har77] (see below). We will conclude this section with some further evidence.
5.2.7.
The Beilinson-Soulé conjecture. Actually the above conjecture in the range $n<1$ is already well-known as the Beilinson-Soulé conjecture, in its “strong” form — this terminology appears in the introduction of [Lev93]. By opposition, the “weak” form of the Beilinson-Soulé conjecture asserts the vanishing of motivic cohomology in degrees $<0$ (see e.g. [SV00, §3, Be2]). [^49] The integral version is equivalent to the rational one: see [Kah05, Lem. 24]. In the sequel, we will simply use the terminology Beilinson-Soulé conjecture for the strong form of the conjecture.